# Comprehensive Evaluation of OCT-based Automated Segmentation of Retinal Layer, Fluid and Hyper-Reflective Foci: Impact on Diabetic Retinopathy Severity Assessment

### v3

@@ V Limitation and Future Work
First, manual labelling imperfections impact both model performance and thickness measurements. Retinal layer segmentation is inherently challenging due to subtle boundary variations and overlapping structures. In cases where excessive fluid penetrates the layer boundaries, some portions of the layer become invisible or physically diminished. More clinical expertise is needed to segment the extreme instances properly. HRF segmentation suffers from inconsistencies in ground-truth annotations, as small, widely distributed foci are challenging to delineate manually. Interestingly, in some cases, automated models provided more precise segmentations than the ground-truth labels. For example, for the first B-scan in Figure 5(a), the predictions of SegFormer and SwinUNETR have better RNFL segmentation than the ground truth with a smoother and more precise layer boundary. For the first B-scan in Figure 5(c), no fluid is manually annotated, but the segmentation models, except for SegFormer, predict potential intra-retinal fluid across the OPL and ONL+IS regions. Additionally, using more pre-processing and post-processing techniques may help improve the performance, such as the pixel-wise relative positional map as an extra input and random forest classifier as a label refiner(Ma et al., 2021).
@@
Second, additional model comparisons may be necessary to provide a more comprehensive evaluation of segmentation approaches. While SwinUNETR and VM-UNet demonstrated superior performance, other architectures excel in certain perspectives. For example, MedSAM enables universal medical image segmentation with zero-shot capabilities(Ma et al., 2023). The self-supervised few-shot semantic segmentation can be used for a limited number of labels(Ouyang et al., 2022). A broader comparison across multiple deep learning models could offer more insights into the trade-offs between performance, efficiency, and generalizability.
@@
Third, this study’s cross-sectional nature limits its ability to track disease progression over time. Longitudinal studies would provide better insights into how retinal layer thickness evolves in DR. For example, several studies report RNFL/GCL thinning during the progression of DR, which has become one of the most important preclinical biomarkers for DR severity evaluation(Bhaskaran et al., 2023)(Oshitari et al., 2009)(Park et al., 2011). Additionally, while the sample size is sufficient to detect significant differences, it may limit the generalizability of the findings. A larger dataset encompassing a broader range of DR severities and treatment histories could provide more robust conclusions. In our study, no significant RNFL thickness difference is found between the NPDR and PDR groups. Fig 7 shows explicit GCL+IPL thinning of PDR in terms of the mean and interquartile range, but only the NI sector exhibits marginal significance($p=0.058$). Expanding the cohort to include more diverse patient populations may help improve the applicability of the findings across different clinical settings.
@@
Lastly, integrating multi-modal imaging techniques such as OCT angiography (OCTA) could provide additional insights into the vascular changes associated with DR. For example, Alam et al. discovered the difference in vascular complexity features between NPDR and PDR patients(Alam et al., 2021). Multiple OCT parameters are significantly correlated with DR severity(Laotaweerungsawat et al., 2020). Combining structural OCT findings with functional vascular imaging may improve disease characterization and facilitate more targeted therapeutic interventions.
@@ VI Conclusion
This study highlights the strengths and limitations of current deep learning-based segmentation models in analyzing diabetic retinopathy (DR)-related structural changes. Both SwinUNETR and VM-UNet exhibit strong performance, particularly in segmenting complex retinal layers and fluid regions. However, segmentation of fluid and HRF remains challenging due to their small size and dispersed distribution. Analysis of retinal layer thickness differences between NPDR and PDR reveals distinct structural alterations, with significant differences observed in the OPL, ONL+IS, fluid, and HRF distributions. The varying relationships between visual acuity and these structural changes in NPDR versus PDR suggest a progression from adaptive retinal remodelling in NPDR to pathological neurodegeneration and edema-driven vision loss in PDR, reinforcing the importance of early detection and intervention.
@@ VIII Declaration of competing interest
Layer Model SI NI II TI SO NO IO TO
$\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr
RNFL GT -0.046 0.526 0.913 -0.044 0.56 0.913 -0.001 0.984 0.988 0.071 0.178 0.913 -0.012 0.837 0.967 -0.011 0.85 0.967 0.006 0.927 0.967 0.111 0.06 0.913
Swin -0.043 0.531 0.913 -0.08 0.197 0.913 -0.02 0.742 0.952 -0.005 0.907 0.967 0.011 0.858 0.967 0.008 0.909 0.967 0.034 0.598 0.913 0.08 0.178 0.913
VM -0.066 0.283 0.913 -0.083 0.146 0.913 -0.017 0.782 0.967 0.023 0.548 0.913 -0.039 0.473 0.913 -0.032 0.531 0.913 0.02 0.737 0.952 0.086 0.094 0.913
GCL+IPL GT -0.03 0.391 0.913 -0.027 0.401 0.913 -0.001 0.988 0.988 -0.044 0.236 0.913 0.022 0.5 0.913 0.034 0.266 0.913 0.016 0.644 0.932 -0.021 0.527 0.913
Swin -0.042 0.272 0.913 -0.008 0.832 0.967 0.014 0.721 0.941 -0.053 0.186 0.913 0.033 0.313 0.913 0.044 0.197 0.913 0.036 0.318 0.913 -0.016 0.655 0.932
VM -0.056 0.153 0.913 -0.035 0.279 0.913 -0.009 0.797 0.967 -0.062 0.124 0.913 0.022 0.519 0.913 -0.048 0.528 0.913 0.006 0.875 0.967 -0.013 0.71 0.937
INL GT 0.007 0.867 0.967 -0.001 0.968 0.988 -0.023 0.519 0.913 -0.076 0.25 0.913 0.041 0.285 0.913 0.014 0.696 0.937 0.018 0.692 0.937 0.023 0.559 0.913
Swin -0.006 0.863 0.967 0.03 0.294 0.913 0.014 0.682 0.937 -0.098 0.155 0.913 0.054 0.16 0.913 0.058 0.08 0.913 0.064 0.091 0.913 0.023 0.563 0.913
VM -0.047 0.237 0.913 0.012 0.709 0.937 -0.025 0.517 0.913 -0.093 0.134 0.913 0.024 0.552 0.913 0.025 0.428 0.913 0.03 0.43 0.913 0.005 0.911 0.967
OPL GT 0.078 0.244 0.913 0.039 0.616 0.913 0.02 0.86 0.967 0.042 0.674 0.936 0.093 0.012 0.913 0.033 0.392 0.913 0.035 0.573 0.913 0.054 0.304 0.913
Swin 0.04 0.467 0.913 0.043 0.405 0.913 0.063 0.27 0.913 0.031 0.666 0.932 0.051 0.226 0.913 0.029 0.349 0.913 0.055 0.212 0.913 0.051 0.252 0.913
VM 0.039 0.605 0.913 -0.01 0.894 0.967 0.07 0.426 0.913 0.043 0.701 0.937 0.05 0.263 0.913 0.022 0.441 0.913 0.037 0.43 0.913 0.065 0.209 0.913
ONL+IS GT -0.158 0.029 0.913 -0.128 0.016 0.913 -0.038 0.501 0.913 -0.122 0.079 0.913 -0.081 0.244 0.913 -0.054 0.311 0.913 -0.025 0.64 0.932 -0.035 0.579 0.913
Swin -0.056 0.359 0.913 -0.045 0.369 0.913 0.005 0.923 0.967 -0.041 0.436 0.913 -0.034 0.55 0.913 -0.078 0.229 0.913 -0.005 0.918 0.967 0.322 0.38 0.913
VM -0.075 0.215 0.913 -0.08 0.124 0.913 -0.019 0.725 0.941 -0.101 0.083 0.913 -0.034 0.573 0.913 -0.008 0.885 0.967 -0.016 0.757 0.96 -0.049 0.435 0.913
EZ GT -0.028 0.597 0.913 0.065 0.555 0.913 0.112 0.125 0.913 0.065 0.65 0.932 -0.007 0.843 0.967 0.011 0.807 0.967 0.02 0.579 0.913 0.022 0.608 0.913
Swin -0.113 0.102 0.913 -0.027 0.575 0.913 0.088 0.144 0.913 -0.079 0.402 0.913 -0.041 0.192 0.913 -0.033 0.244 0.913 -0.03 0.219 0.913 0.077 0.29 0.913
VM -0.053 0.149 0.913 0.026 0.619 0.913 0.099 0.078 0.913 -0.018 0.85 0.967 -0.024 0.412 0.913 -0.026 0.394 0.913 -0.014 0.571 0.913 0.013 0.709 0.937
OS GT -0.027 0.564 0.913 0.001 0.98 0.988 0.005 0.912 0.967 -0.01 0.837 0.967 -0.012 0.76 0.96 -0.003 0.943 0.975 0.009 0.821 0.967 0.012 0.746 0.952
Swin 0.019 0.585 0.913 0.007 0.85 0.967 0.018 0.586 0.913 0.023 0.509 0.913 -0.004 0.884 0.967 -0.129 0.328 0.913 0.003 0.919 0.967 0.027 0.453 0.913
VM 0.013 0.72 0.941 0.008 0.813 0.967 -0.001 0.984 0.988 0.008 0.831 0.967 -0.001 0.96 0.985 -0.014 0.683 0.937 0.004 0.894 0.967 0.064 0.324 0.913
RPE GT 0.007 0.822 0.967 -0.019 0.57 0.913 0.006 0.834 0.967 0.005 0.889 0.967 0.041 0.137 0.913 0.013 0.661 0.932 0.025 0.344 0.913 -0.012 0.66 0.932
Swin 0.009 0.787 0.967 -0.003 0.933 0.97 0.023 0.426 0.913 -0.011 0.687 0.937 0.024 0.374 0.913 0.029 0.395 0.913 0.021 0.43 0.913 -0.029 0.192 0.913
VM 0.294 0.377 0.913 0.018 0.528 0.913 0.026 0.279 0.913 1.302 0.101 0.913 0.003 0.884 0.967 0.114 0.118 0.913 0.011 0.602 0.913 1.525 0.072 0.913
Fluid GT 0.298 0.441 0.913 0.595 0.111 0.913 0.195 0.62 0.913 -0.175 0.607 0.913 0.857 0.01 0.913 0.062 0.921 0.967 -0.123 0.832 0.967 0.1 0.8 0.967
Swin 0.01 0.977 0.988 0.355 0.275 0.913 0.019 0.954 0.983 -0.159 0.631 0.923 0.191 0.668 0.932 -0.241 0.562 0.913 -0.551 0.215 0.913 0.243 0.446 0.913
VM 0.213 0.564 0.913 0.687 0.055 0.913 0.194 0.59 0.913 -0.094 0.779 0.967 0.291 0.375 0.913 0.321 0.6 0.913 0.376 0.377 0.913 0.399 0.239 0.913
HRF GT 0.351 0.468 0.913 0.93 0.04 0.913 0.377 0.273 0.913 0.583 0.158 0.913 0.248 0.393 0.913 0.32 0.351 0.913 0.35 0.381 0.913 0.55 0.099 0.913
Swin 0.068 0.874 0.967 0.886 0.056 0.913 0.273 0.398 0.913 0.389 0.226 0.913 0.228 0.422 0.913 0.204 0.463 0.913 0.362 0.313 0.913 0.536 0.067 0.913
VM -0.275 0.519 0.913 0.85 0.084 0.913 0.279 0.366 0.913 0.377 0.329 0.913 0.233 0.418 0.913 0.333 0.281 0.913 0.286 0.415 0.913 0.601 0.077 0.913

### v4

@@ V Limitation and Future Work
First, manual labelling imperfections impact both model performance and thickness measurements. Retinal layer segmentation is inherently challenging due to subtle boundary variations and overlapping structures. In cases where excessive fluid penetrates the layer boundaries, some portions of the layer become invisible or physically diminished. More clinical expertise is needed to segment the extreme instances properly. HRF segmentation suffers from inconsistencies in ground-truth annotations, as small, widely distributed foci are challenging to delineate manually. Interestingly, in some cases, automated models provided more precise segmentations than the ground-truth labels. For example, for the first B-scan in Figure 5(a), the predictions of SegFormer and SwinUNETR have better RNFL segmentation than the ground truth, with a smoother and more precise layer boundary. For the first B-scan in Figure 5(c), no fluid is manually annotated, but the segmentation models, except for SegFormer, predict potential intra-retinal fluid across the OPL and ONL+IS regions. Additionally, using more pre-processing and post-processing techniques may help improve the performance, such as the pixel-wise relative positional map as an extra input and a random forest classifier as a label refiner[19].
@@
Second, additional model comparisons may be necessary to provide a more comprehensive evaluation of segmentation approaches. While SwinUNETR and VM-UNet demonstrated superior performance, other architectures excel in certain perspectives. For example, MedSAM enables universal medical image segmentation with zero-shot capabilities[35]. Zhu et al. propose MedSAM-2 to perform medical segmentation tasks as a video object tracking problem[36]. The self-supervised few-shot semantic segmentation can be used for a limited number of labels[37]. However, many of these approaches rely on real-time user input or manual corrections, which diverges from the goal of this study—to assess whether widely adopted deep learning models can autonomously produce clinically reliable results. A broader comparison across multiple deep learning models could offer more insights into the trade-offs between performance, efficiency, and generalizability. The proposed model should also be validated from an external dataset, especially acquired from multiple OCT devices. To the best of our knowledge, there are currently no public datasets available with a comprehensive layer and fluid segmentation from severe DR patients. To advance benchmarking and promote the development of robust and generalizable segmentation models, greater efforts from the research community are needed to share diverse, well-annotated datasets.
@@
Third, this study’s cross-sectional nature limits its ability to track disease progression over time. Longitudinal studies would provide better insights into how retinal layer thickness evolves in DR. For example, several studies report RNFL/GCL thinning during the progression of DR, which has become one of the most important preclinical biomarkers for DR severity evaluation[38, 39, 40]. Additionally, while the sample size is sufficient to detect significant differences, it may limit the generalizability of the findings. A larger dataset encompassing a broader range of DR severities and treatment histories could provide more robust conclusions. In our study, no significant RNFL thickness difference is found between the NPDR and PDR groups. Fig 7 shows explicit GCL+IPL thinning of PDR in terms of the mean and interquartile range, but only the NI sector exhibits marginal significance($p=0.058$). Expanding the cohort to include more diverse patient populations may help improve the applicability of the findings across different clinical settings.
@@
Lastly, integrating multi-modal imaging techniques such as OCT angiography (OCTA) could provide additional insights into the vascular changes associated with DR. For example, Alam et al. discovered the difference in vascular complexity features between NPDR and PDR patients[41]. Multiple OCT parameters are significantly correlated with DR severity[42]. Combining structural OCT findings with functional vascular imaging may improve disease characterization and facilitate more targeted therapeutic interventions.
@@ VI Conclusion
This study highlights the strengths and limitations of current deep learning-based segmentation models in analyzing diabetic retinopathy (DR)-related structural changes. Both SwinUNETR and VM-UNet exhibit strong performance, particularly in segmenting complex retinal layers and fluid regions. However, segmentation of fluid and HRF remains challenging due to their small size and dispersed distribution. Analysis of retinal layer thickness differences between NPDR and PDR reveals distinct structural alterations, with significant differences observed in the RNFL, GCL+IPL, OS, and fluid. The varying relationships between visual acuity and these structural changes in NPDR versus PDR suggest a progression from adaptive retinal remodelling in NPDR to pathological neurodegeneration and edema-driven vision loss in PDR, reinforcing the importance of early detection and intervention.
@@ VIII Declaration of competing interest
IX Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author(s) used ChatGPT (OpenAI) in order to improve the language and clarity of the manuscript. After using this tool, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
@@ IX Declaration of generative AI and AI-assisted technologies in the writing process
C SI NI II TI SO NO IO TO
$\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR
RNFL GT / / / 0.093 0.192 0.485 0.115 0.110 0.393 0.035 0.596 0.838 0.117 0.028 0.263 0.110 0.047 0.279 0.120 0.022 0.222 0.020 0.770 0.889 0.130 0.034 0.264
Swin / / / 0.075 0.263 0.588 0.071 0.251 0.575 -0.004 0.954 0.949 0.009 0.847 0.923 0.101 0.079 0.345 0.075 0.280 0.601 -0.045 0.483 0.784 0.088 0.145 0.441
VM / / / 0.127 0.031 0.263 0.120 0.032 0.263 0.049 0.416 0.732 0.095 0.013 0.197 0.111 0.034 0.264 0.099 0.039 0.277 -0.000 0.995 0.960 0.135 0.008 0.156
GCL+IPL GT / / / -0.070 0.045 0.277 -0.052 0.111 0.393 -0.089 0.010 0.177 -0.107 0.004 0.096 -0.005 0.877 0.923 -0.025 0.408 0.729 -0.059 0.076 0.345 -0.047 0.178 0.465
Swin / / / -0.059 0.122 0.393 -0.059 0.117 0.393 -0.092 0.017 0.206 -0.094 0.017 0.206 -0.010 0.773 0.889 -0.031 0.362 0.677 -0.066 0.058 0.326 -0.031 0.378 0.692
VM / / / -0.054 0.172 0.460 -0.061 0.065 0.328 -0.113 0.002 0.052 -0.095 0.018 0.206 0.001 0.979 0.960 0.094 0.169 0.460 -0.053 0.115 0.393 -0.041 0.254 0.577
INL GT / / / 0.003 0.939 0.939 0.003 0.921 0.932 0.006 0.865 0.923 0.088 0.173 0.460 0.008 0.822 0.923 -0.024 0.488 0.786 -0.050 0.264 0.588 0.000 0.995 0.960
Swin / / / -0.047 0.198 0.494 -0.023 0.439 0.742 0.006 0.849 0.923 0.060 0.374 0.692 -0.011 0.771 0.889 -0.012 0.722 0.870 -0.058 0.116 0.393 0.014 0.725 0.870
VM / / / -0.052 0.208 0.500 -0.012 0.717 0.870 0.017 0.667 0.870 0.078 0.219 0.508 -0.013 0.728 0.870 -0.016 0.594 0.838 -0.041 0.269 0.592 0.025 0.536 0.817
OPL GT / / / 0.011 0.884 0.925 0.052 0.511 0.802 -0.016 0.878 0.923 -0.006 0.956 0.949 0.024 0.537 0.817 -0.025 0.519 0.804 -0.063 0.306 0.641 0.050 0.358 0.676
Swin / / / 0.029 0.595 0.838 0.058 0.279 0.601 0.035 0.555 0.824 0.034 0.641 0.863 -0.005 0.903 0.932 -0.021 0.498 0.792 -0.062 0.145 0.441 0.017 0.723 0.870
VM / / / -0.019 0.813 0.923 0.033 0.670 0.870 -0.001 0.993 0.960 0.000 0.997 0.960 -0.028 0.533 0.817 -0.028 0.319 0.651 -0.080 0.084 0.349 0.020 0.713 0.870
ONL+IS GT / / / -0.103 0.156 0.456 -0.046 0.430 0.738 -0.107 0.066 0.328 -0.096 0.165 0.460 -0.122 0.067 0.328 -0.021 0.707 0.870 -0.097 0.065 0.328 -0.034 0.605 0.838
Swin / / / -0.113 0.068 0.328 -0.026 0.630 0.862 -0.101 0.079 0.345 -0.124 0.020 0.214 -0.088 0.111 0.393 0.056 0.403 0.726 -0.108 0.027 0.263 0.292 0.316 0.650
VM / / / -0.079 0.204 0.495 -0.005 0.934 0.937 -0.069 0.220 0.508 -0.115 0.059 0.326 -0.090 0.131 0.416 0.007 0.909 0.932 -0.095 0.051 0.298 -0.059 0.378 0.692
EZ GT -0.296 0.185 0.477 -0.008 0.875 0.923 -0.063 0.544 0.822 -0.013 0.867 0.923 -0.061 0.676 0.870 0.013 0.704 0.870 0.010 0.816 0.923 0.004 0.904 0.932 0.025 0.569 0.828
Swin -0.248 0.098 0.375 -0.053 0.450 0.750 -0.008 0.873 0.923 -0.043 0.491 0.786 -0.136 0.187 0.478 -0.010 0.748 0.881 0.012 0.671 0.870 -0.018 0.459 0.754 0.057 0.429 0.738
VM -0.406 0.014 0.206 -0.013 0.741 0.881 -0.038 0.482 0.784 -0.075 0.202 0.495 -0.144 0.152 0.453 -0.023 0.428 0.738 -0.001 0.983 0.960 -0.024 0.355 0.676 -0.007 0.834 0.923
OS GT -0.143 0.098 0.375 -0.125 0.006 0.124 -0.076 0.083 0.349 -0.111 0.011 0.181 -0.065 0.172 0.460 -0.131 0.000 0.008 -0.103 0.002 0.059 -0.155 0.000 0.001 -0.115 0.001 0.029
Swin -0.025 0.603 0.838 -0.070 0.045 0.277 -0.009 0.815 0.923 -0.055 0.095 0.375 -0.055 0.117 0.393 -0.110 0.000 0.002 0.083 0.516 0.804 -0.126 0.000 0.002 -0.111 0.001 0.029
VM -0.030 0.568 0.828 -0.076 0.031 0.263 -0.056 0.121 0.393 -0.063 0.068 0.328 -0.049 0.165 0.460 -0.095 0.000 0.020 -0.070 0.041 0.277 -0.134 0.000 0.000 -0.037 0.550 0.822
RPE GT 0.052 0.219 0.508 0.007 0.830 0.923 0.031 0.351 0.676 -0.011 0.719 0.870 0.019 0.568 0.828 0.004 0.871 0.923 -0.003 0.925 0.932 -0.026 0.336 0.664 -0.009 0.745 0.881
Swin 0.023 0.571 0.828 0.004 0.895 0.932 -0.016 0.654 0.870 -0.009 0.773 0.889 0.010 0.706 0.870 0.011 0.686 0.870 -0.032 0.359 0.676 -0.046 0.084 0.349 -0.022 0.328 0.660
VM -0.262 0.599 0.838 -0.154 0.633 0.862 -0.015 0.596 0.838 -0.025 0.300 0.633 -0.665 0.412 0.731 -0.004 0.850 0.923 -0.082 0.271 0.592 -0.038 0.060 0.326 -0.808 0.355 0.676
Fluid GT 0.333 0.113 0.393 0.345 0.087 0.355 0.519 0.019 0.206 0.292 0.155 0.456 0.278 0.146 0.441 0.029 0.866 0.923 0.003 0.989 0.960 -0.226 0.204 0.495 0.344 0.044 0.277
Swin 0.398 0.168 0.460 0.681 0.030 0.263 0.768 0.017 0.206 0.476 0.159 0.458 0.543 0.096 0.375 0.150 0.702 0.870 0.316 0.432 0.738 0.766 0.076 0.345 0.168 0.639 0.863
VM 0.280 0.330 0.660 0.453 0.177 0.465 0.734 0.042 0.277 0.355 0.332 0.660 0.488 0.143 0.441 -0.360 0.299 0.633 -0.269 0.610 0.840 -0.417 0.315 0.650 0.163 0.665 0.870
HRF GT 0.037 0.836 0.923 0.051 0.725 0.870 0.096 0.548 0.822 -0.041 0.767 0.889 0.013 0.924 0.932 0.046 0.690 0.870 0.087 0.434 0.738 0.057 0.650 0.870 0.070 0.502 0.794
Swin -0.363 0.384 0.697 -0.344 0.453 0.750 0.053 0.912 0.932 -0.572 0.070 0.328 0.031 0.922 0.932 -0.611 0.039 0.277 -0.143 0.582 0.838 -0.442 0.218 0.508 0.047 0.865 0.923
VM 0.173 0.720 0.870 -0.355 0.445 0.746 -0.000 0.999 0.960 -0.490 0.121 0.393 0.125 0.724 0.870 -0.617 0.043 0.277 -0.273 0.346 0.676 -0.544 0.122 0.393 0.050 0.877 0.923